f(x)= - (x IxI )/ (1+x2), then f-1(x) equals? Why?

Grade:12th Pass

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello
when x is +
y=-x^2/(1+x^2)
=>y+1=1/1+x^2
=>x^2=-y/1+y
similarly when x is -
x^2=y/1-y
Thanks & Regards
Arun Kumar
Btech IIT delhi
Askiitians Faculty